OpenStudy (anonymous):

How do I find the zeroes of this function: f(x)=x^2+6x-1

OpenStudy (seratul):

Isolate the variable. Move x to one side. I'll help you start. -1 in both sides.

OpenStudy (seratul):

I mean +1

OpenStudy (anonymous):

Ok, so it would be 1=x^2+6x.

OpenStudy (snowsurf):

T find the zeros you need to find the factors that will give you zeros. x^2+6x-1 = 0 However, by looking at this it does not look it will factor nicely so you using the quadratic formula will be your best bet.

OpenStudy (anonymous):

What is the quadratic formula?

OpenStudy (snowsurf):

\[x = \frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\] Apply this quadratic formula to find the values that will equal zero.

OpenStudy (snowsurf):

ax^2 +bx + c =0

OpenStudy (anonymous):

Isn't there an easier way besides doing the formula?

OpenStudy (snowsurf):

Did you go over quadratic formula in class?

OpenStudy (anonymous):

No. We learned a different way. I just cant tell what I did from my notes.

OpenStudy (snowsurf):

There are different ways solve this you can use quadratic, you can do factoring, you can do completing the square. Review your notes and see what way you were taught from class.

OpenStudy (anonymous):

Well, we did factoring last chapter, so I am pretty sure it was by factoring. We are also doing completing the square this chapter, but I just want to do whichever one is easier.

OpenStudy (snowsurf):

Let's try factoring By looking at the x^2 term we can only have x*x and by looking at the -1 at the end we can only have (1)(-1) = -1 But we need an way to sum the middle term to 6. From these only factors that cannot be done. That is why this cannot be factored. So completing the square would be our option. So try showing your steps to completing the square. You can attach a photo of your work.